OpenStudy (kamille):

I need the to find smallest positive solution \[(\sin \pi x)(cosx-2)=0\] \[\sin \pi x=0\] or \[cosx - 2=0\] so equation of \[cosx - 2=0\] doesn't have a solution,does it?

OpenStudy (anonymous):

Correct, cos(x) - 2 = 0 doesn't have a solution. Now if the sine of your particular angle pi*x = 0, then what must the angle be?

OpenStudy (kamille):

or do i need to write like \[cosx-2=0\] \[x-2=\pm \arccos a +2pik\]

OpenStudy (anonymous):

0 isn't a positive solution...

OpenStudy (kamille):

pi*x=0|:pi x=0?

OpenStudy (anonymous):

@Kamille Did you mean \[\large \cos(x-2)=0\] instead?

OpenStudy (anonymous):

Because that wouldn't work.

OpenStudy (kamille):

iknow i need to sinx=0 x=pik

OpenStudy (kamille):

no, I meant cosx−2=0

OpenStudy (anonymous):

If sin(pi*x) = 0, then pi*x must equal 0, pi, or some multiple of pi. Hence, if you want a positive solution, x = 1. Any other solution will either be negative, zero, or greater than 1.

OpenStudy (kamille):

oh, okay, thank you!